OpenStudy (anonymous):

Sarah and Billy built their own tree houses that were the same size. It took Billy 5 hours to build his treehouse and Sarah took 8 hours. How much time would it take for Sarah and Billy built a treehouse together? Write a rational expression based on the word problem.

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

I do these like parallel resistor problems in physics.\[\Large \frac{ 1 }{ t }=\frac{ 1 }{ 5 }+\frac{ 1 }{8 }\]Solve for t.

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

(1/5 + 1/8)(t) = 1

OpenStudy (anonymous):

cross multiply correct?

OpenStudy (radar):

A tip or shortcut, when two are involved, use product over sum: (5)(8)/(5 + 8) or 40/13

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

could someone walk me through how to solve for t?

OpenStudy (anonymous):

t = 1/ (1/5 + 1/8)

OpenStudy (anonymous):

what? can you use the equation editor

OpenStudy (anonymous):

\[t= \frac{1 }{ 1/5 + 1/8 }\]

OpenStudy (anonymous):

hmm next step?

OpenStudy (anonymous):

@radar @RadEn

OpenStudy (radar):

I use product over sum 40/13, but if you insist on using the old fashion way, the common denominator would be 40